A first order reaction has a rate constant 1.15x10-3 s-1. How long will 5 gram o this reactant take reduce to 3 gram
Answer #1

we have:

[A]o = 5 g

[A] = 3 g

k = 1.15*10^-3 s-1

use integrated rate law for 1st order reaction

ln[A] = ln[A]o - k*t

ln(3) = ln(5) - 1.15*10^-3*t

1.099 = 1.609 - 1.15*10^-3*t

1.15*10^-3*t = 0.5108

t = 4.442*10^2 s

Answer: 4.44*10^2 s